OpenStudy (anonymous):

Which of the following is an equation for a function that has the shape of , but is shifted right 4 units and up one unit? A. f(x)=(x+4)^3+1 B. f(x)=(x+1)^3-4 c.f(x)=(x-4)^3+1 d.f(x)=(x+1)^3+4 Please & Thank YOU

OpenStudy (anonymous):

C!

OpenStudy (anonymous):

c.f(x)=(x-4)^3+1

OpenStudy (anonymous):

if I hav the equation f(x) = (x+a)^3 + b a= is horizontal shift, where - a is right and +a is left b = vertical shift, where + b is up, and -b is down